OpenStudy (camijacobson):

Jerry's bowling score was 230 last week, but only 190 this week. What was the percent decrease in Jerry's score? A. 2.11% B. 1.74% C. 17.4% D. 21.1%

OpenStudy (anonymous):

C because new - original / original, So 190 - 230/230 = -40/230 = 17.39.... so 17.4%

OpenStudy (mathstudent55):

To find a percent change do this: 1. Subtract the new number minus the old number. 2. Divide the difference by the old number. 3. Multiply by 100. If the answer is positive, it is a percent increase. If the answer is negative, it is a percent decrease.

OpenStudy (anonymous):

Need anymore help?

OpenStudy (camijacobson):

not at the moment, but I'll let you know if i do! @wally.dwyer

OpenStudy (anonymous):

if you need anything I will be there